OpenStudy (anonymous):

solve the indefinite integral ∫ (dx)/(√1-(x/3)^2)

OpenStudy (anonymous):

\[\int\frac{dx}{\sqrt{1-\frac{x^2}{9}}}~~?\]

OpenStudy (anonymous):

If that is your integral, substitute \(x=3\sin u\), then \(dx=3\cos u~du\): \[\int\frac{3\cos u}{\sqrt{1-\frac{1}{9}\left(3\sin u\right)^2}}~du\\ \int\frac{3\cos u}{\sqrt{1-\sin^2u}}~du\\ \int\frac{3\cos u}{\cos u}~du\\ 3\int du\]